Simon Company’s year-end balance sheets follow.

At December 31 Current Yr 1 Yr Ago 2 Yrs Ago
Assets
Cash $ 32,000 $ 38,160 $ 40,951
Accounts receivable, net 91,808 68,103 53,011
Merchandise inventory 118,965 89,993 55,887
Prepaid expenses 10,201 10,116 4,462
Plant assets, net 294,873 265,910 247,289
Total assets $ 547,847 $ 472,282 $ 401,600
Liabilities and Equity
Accounts payable $ 139,142 $ 79,816 $ 52,481
Long-term notes payable secured by
mortgages on plant assets
100,935 107,539 90,529
Common stock, $10 par value 162,500 162,500 162,500
Retained earnings 145,270 122,427 96,090
Total liabilities and equity $ 547,847 $ 472,282 $ 401,600

(1-a) Compute the current ratio for each of the three years.
(1-b) Did the current ratio improve or worsen over the three year period?
(2-a) Compute the acid-test ratio for each of the three years.
(2-b) Did the acid-test ratio improve or worsen over the three year period?

Compute the current ratio for each of the three years.

Current Ratio
Choose Numerator: / Choose Denominator: = Current Ratio
/ = Current ratio
Current Yr: / = to 1
1 Yr Ago: / = to 1
2 Yrs Ago: / = to 1

Did the current ratio improve or worsen over the three year period?

Current ratio

Compute the acid-test ratio for each of the three years.

Acid-test ratio
Choose Numerator: / Choose Denominator: = Acid-Test Ratio
+ + / = Acid-test ratio
Current Yr: + + / = to 1
1 Yr Ago: + + / = to 1
2 Yrs Ago: + + / = to 1

Did the acid-test ratio improve or worsen over the three year period?

Acid-test ratio
Answer #1

Compute the current ratio for each of the three years.

Current Ratio
Choose Numerator: / Choose Denominator: = Current Ratio
Current assets / Current liabilities = Current ratio
Current Yr: 252974 / 139142 = 1.82 to 1
1 Yr Ago: 206372 / 79816 = 2.59 to 1
2 Yrs Ago: 154311 / 52481 = 2.94 to 1

Did the current ratio improve or worsen over the three year period?

Current ratio Worsen

Compute the acid-test ratio for each of the three years.

Acid-test ratio
Choose Numerator: / Choose Denominator: = Acid-Test Ratio
Cash + Account receivable + Short term investment / Current liabilities = Acid-test ratio
Current Yr: 32000 + 91808 + 0 / 139142 = 0.89 to 1
1 Yr Ago: 38160 + 68103 + 0 / 79816 = 1.33 to 1
2 Yrs Ago: 40951 + 53011 + 0 / 52481 = 1.79 to 1

Did the acid-test ratio improve or worsen over the three year period?

Acid test ratio is worsen